OpenStudy (anonymous):

a linear regression was conducted on the following data points. A(1, 11) B(3, 7) C(5, 2) D(7, 1) E(9, -2) The residual for which point has the largest absolute value?

OpenStudy (amistre64):

looks like you either need to develop the linear regression equation, or use the one they give you if they gave you one, to determine the difference between a stated y, and the lr y value

OpenStudy (anonymous):

@ashleybvb

OpenStudy (amistre64):

\[Y=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}(X-\bar X)+\bar Y\]

OpenStudy (anonymous):

I don't understand that equation because I cant afford the calculator

OpenStudy (amistre64):

its 5 points ... which can be pretty much done by hand i spose; if organised in the right way

OpenStudy (amistre64):

x*x x y x*y 1^2 (1, 11) 1(11) 3^2 (3, 7) 3(7) 5^2 (5, 2) 5(2) 7^2 (7, 1) 7(1) 9^2 (9, -2) 9(-2) --------------------- totals of columns

OpenStudy (amistre64):

when we know the totals of each column, we can then construct the linear regression equation from them

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

so you basically multiply each one and the one that is the highest is the answer?

OpenStudy (amistre64):

if they do not give you an equation to work with, then you need to construct one from points

OpenStudy (amistre64):

there is no "simple" way to go about it in my view

OpenStudy (amistre64):

x*x x y x*y 1^2 (1, 11) 1(11) 3^2 (3, 7) 3(7) 5^2 (5, 2) 5(2) 7^2 (7, 1) 7(1) 9^2 (9, -2) 9(-2) --------------------- 165 25 19 31 n = 5 points, avgX = 25/5 = 5 avgY = 19/5 \[Y=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}(X-\bar X)+\bar Y\] \[Y=\frac{5(31)-(25)(19)}{5(165)-(25)(25)}(X-5)+\frac{19}{5}\]i believe would be the regression equation

OpenStudy (amistre64):

to find the difference between Y and a given y point, let X=x and subtract y from both sides \[Y-y=\frac{5(31)-(25)(19)}{5(165)-(25)(25)}(x-5)+\frac{19}{5}-y\] and test out each x,y point

OpenStudy (amistre64):

10.2 - 11 .8 7 - 7 0 3.8 - 2 1.8 .6 - 1 .4 -2.6 + 2 .6

OpenStudy (anonymous):

@amistre64 so the answer would be c(5, 2)

OpenStudy (amistre64):

assuming i didnt mess it up along the way, i would say C is the one with the furthest gap, yes

OpenStudy (anonymous):

thank you